OpenStudy (anonymous):

Runner A crosses the starting line of a marathon and runs at an average pace of 5.6 miles per hour. Half an hour later, Runner B crosses the starting line and runs at an average rate of 6.4 miles per hour. If the length of the marathon is 26.2 miles, which runner will finish ahead of the other? Explain.

OpenStudy (anonymous):

Runner A; Runner B will not be able to catch Runner A in the time it takes Runner A to complete the 26.2 mile course. Runner B; Runner B will catch up to runner A 3.5 hours after Runner A crosses the starting line. Runner B; Runner B will pass Runner A and finish more than half an hour ahead of Runner A. Runner B; Runner B will catch up to Runner A 4 hours after Runner A crosses the starting line.

OpenStudy (anonymous):

@Deadly_Roses Help?

OpenStudy (deadly_roses):

Runner A: Time it takes for him to finish: t = 26.2/5.6 = 4.679 hours (to complete race) . Runner B: Time it takes for him to finish: t = 26.2/6.4 = 4.094 hours BUT wait, we have to account for the extra .5 hours -- the time it took for him to cross the start line. t = 4.094 + .5 = 4.594 hours . Since Runner B's time is less than Runner A's, we must conclude Runner B will finish first.

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

or C.?

OpenStudy (deadly_roses):

yep runner B will finish first

OpenStudy (deadly_roses):

i think c would be correct but b could be too

OpenStudy (anonymous):

thank you :)

OpenStudy (deadly_roses):

np